# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 2.2 The Singular Value Decomposition
where $r=\min(m,n)$, $\bm{\mathsf{U}}$ and $\bm{\mathsf{V}}$ are orthonormal, and $\mathbf{\Sigma}$ is diagonal. The columns $\{\bm{\mathsf{u}}_{i}\}_{i=1}^{r}$ and $\{\bm{\mathsf{v}}_{i}\}_{i=1}^{r}$ of $\bm{\mathsf{U}}$ and $\bm{\mathsf{V}}$, respectively, are called the left and right singular vectors of $\bm{\mathsf{A}}$. The diagonal elements $\{\sigma_{i}\}_{i=1}^{r}$ of $\mathbf{\Sigma}$ are the singular values of $\bm{\mathsf{A}}$, ordered so that $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$. The rank of $\bm{\mathsf{A}}$ is equal to the number of nonzero singular values.
Suppose that the rank of $\bm{\mathsf{A}}$ is at least $k$. To obtain a rank-$k$ approximation of $\bm{\mathsf{A}}$, we can truncate its SVD after the first $k$ terms, defining $\bm{\mathsf{A}}_{k}=\sum_{i=1}^{k}\sigma_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{v}}_{i}^{*}$. By the Eckart-Young theorem eckart1936, $\bm{\mathsf{A}}_{k}$ is the best possible rank-$k$ approximation of $\bm{\mathsf{A}}$, with approximation error given by the singular values:
@@
We adopt this notation when referring to the computation of the SVD, as well as $\texttt{svd}(\bm{\mathsf{A}},k)$ to represent the function outputting the truncation $\bm{\mathsf{A}}_{k}$. Golub-Reinsch and related algorithms rely on QR decompositions, covered next.
@@ 2.3 The QR Decomposition
Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits a QR decomposition given by
$\underset{m\times n}{\hskip 5.69054pt\bm{\mathsf{A}}\hskip 5.69054pt}=\underset{m\times r}{\bm{\mathsf{Q}}}\underset{r\times n}{\hskip 5.69054pt\bm{\mathsf{R}}\hskip 5.69054pt},$ (3)
where $r=\min(m,n)$, $\bm{\mathsf{Q}}$ is orthonormal and $\bm{\mathsf{R}}$ is upper triangular.
Using a greedy algorithm that builds the QR factorization incrementally (e.g., with Householder reflections Householder1958; cf. Sect 2.4), we can halt the factorization after computation of the first $k$ terms to obtain a “partial” QR factorization:
$\underset{m\times n}{\bm{\mathsf{A}}}\approx\underset{m\times k}{\bm{\mathsf{Q}}_{k}}\underset{k\times n}{\hskip 5.69054pt\bm{\mathsf{R}}_{k}\hskip 5.69054pt}.$ (4)
We let
$\begin{bmatrix}{\bm{\mathsf{Q}},\bm{\mathsf{R}}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \begin{bmatrix}{\bm{\mathsf{Q}}_{k},\bm{\mathsf{R}}_{k}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}},k),$ (5)
denote functions that yield either the full or partial QR factorization of $\bm{\mathsf{A}}$, respectively. Occasionally, if we need only the factor $\bm{\mathsf{Q}}$, whose columns form an orthonormal basis for the range of $\bm{\mathsf{A}}$, we will use the notation
@@ 2.4 The Interpolative and CUR Decompositions
where
$\bm{\mathsf{R}}=\bm{\mathsf{A}}(I,:)$
for $k$ linearly independent rows indexed by $I$;
$\bm{\mathsf{C}}=\bm{\mathsf{A}}(:,J)$
for $k$ linearly independent columns indexed by $J$; and $\bm{\mathsf{U}}=\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$, though in practice, $\bm{\mathsf{U}}$ is often approximated by $\bm{\mathsf{A}}(I,J)^{\dagger}$. The columns of $\bm{\mathsf{C}}$ or rows of $\bm{\mathsf{R}}$ are known as skeletons in an ID or CUR decomposition.
@@
where the coefficient $\sqrt{k+1}$ is optimal and cannot be improved; here, $\bm{\mathsf{A}}_{k}$ is the optimal rank-$k$ approximation given by the truncated SVD.
In practice, greedy pivoting algorithms are frequently used to choose skeletons for ID and CUR decompositions. We discuss several popular choices in the following sections, including a recently proposed algorithm that achieves (12) with the best known complexity.
QR with Column Pivoting (CPQR)
Column-pivoted QR (CPQR) is the algorithm behind the qr function from Sect. 2.3. In particular, column pivoting may be required for numerical stability, and the outputs of qr($\bm{\mathsf{A}}$) = $\bm{\mathsf{Q}},\bm{\mathsf{R}},P]$ represent the factorization $\bm{\mathsf{A}}\bm{\mathsf{P}}=\bm{\mathsf{Q}}\bm{\mathsf{R}}$ for some permutation matrix $\bm{\mathsf{P}}\in^{n\times n}$ such that $\bm{\mathsf{P}}=\bm{\mathsf{I}}(:,P)$. This factorization can be achieved through a sequence of permutations and Householder reflections that rank-one update the “active submatrix” of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ by greedy pivoting, based on squared column norms.
Assuming $m\geq n$, let $\bm{\mathsf{A}}^{(t)}$ represent the resulting matrix after $t$ steps of pivoting and updating for $t=0,\ldots,n-2$ (with $\bm{\mathsf{A}}^{(0)}=\bm{\mathsf{A}}$). At the $(t+1)$-th step, CPQR determines the column pivot $j_{t+1}$ by finding the column of maximal Euclidean norm in the active submatrix,
@@
In practice, CPQR chooses pivots in an order that reveals the rank of $\bm{\mathsf{A}}$, so that the chosen skeletons form a good basis for its row or column space. While there are adversarial cases for which CPQR can fail to be rank-revealing (e.g., Kahan-type matrices kahan1966; peters1975), these are rarely encountered, cf. trefethen1990. There are variations of CPQR with better theoretical guarantees, such as rank-revealing QR Chan1987; Ilse1994 or strong rank-revealing QR GuEis1996, but they are more costly to implement.
LU with Partial Pivoting
The LU algorithm with row-wise partial pivoting (LUPP) yields a factorization
$\underset{m\times m}{\hskip 11.38109pt\bm{\mathsf{P}}\hskip 11.38109pt}\underset{m\times n}{\bm{\mathsf{A}}}=\underset{m\times n}{\bm{\mathsf{L}}}\underset{n\times n}{\hskip 5.69054pt\bm{\mathsf{U}}\hskip 5.69054pt},$ (13)
where $\bm{\mathsf{L}}$ is lower triangular with main diagonal entries equal to 1 and off-diagonal entries bounded in magnitude by 1, and $\bm{\mathsf{U}}$ is upper triangular. The function $\bm{\mathsf{L}},\bm{\mathsf{U}},\bm{\mathsf{P}}]=\texttt{lu}(\bm{\mathsf{A}})$ represents the factorization (13) obtained from row-wise partial pivoting.
More specifically, to compute an LU factorization, row pivots are selected from the active submatrix of $\bm{\mathsf{A}}$ that is updated via Schur complements (e.g., (golub2013, Algorithm 3.2.1) (trefethen1997, Algorithm 21.1)). Let $\bm{\mathsf{A}}^{(t)}$ be the resulting matrix after $t$ steps of pivoting and updating for $t=0,1,\ldots,n-2$. At the ($t+1$)-st step of LUPP, the largest-magnitude column entry of the active submatrix is selected as the next pivot entry:
@@
where $\bm{\mathsf{A}}^{(t)}(t+1,t+1)$ has been updated with $\bm{\mathsf{A}}^{(t)}(i_{t+1},t+1)$ following row permutation.
This LUPP algorithm leads to an exponential upper bound on the entries of $\bm{\mathsf{U}}$, which is tight for certain matrices (e.g., Kahan-type matrices kahan1966; peters1975; sorensen2014). Thus, LUPP can fail to be rank-revealing, but such adversarial inputs are rarely encountered in practice trefethen1990. There exist more sophisticated pivoting strategies to lend rank-revealing properties to the LU factorization Chan84; Pan2000; Zhao2005 with better error bounds, but at significantly greater computational cost. Fortunately, we will see in Sect. 4.3 that randomization can equip LUPP with rank-revealing properties comparable to CPQR even in adversarial cases.
Truncated SVD for Near-Optimal Skeletons
@@
By the upper triangularity of $\widehat{\bm{\mathsf{V}}}^{*}$ and the relation (16),
@@
If we add this up over all previous steps, we have
@@
Letting $i=k$ in (20), we achieve the desired bound in (12) in $O(mnk)$ operations. Full details can be found in Osinsky2025.
@@ 3 Randomized Dimension Reduction: Sketching and Sampling
Randomized dimension reduction is frequently associated with randomized embeddings, linear maps drawn from random matrix distributions that satisfy the following condition with high probability. Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be a matrix of rank $k\leq\min(m,n)$ and $\varepsilon\in(0,1)$ a distortion parameter. A linear map $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an embedding of $\bm{\mathsf{A}}$ with distortion $\varepsilon$ if
@@
Usually, $k\leq d\ll m,$ so that the application of $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ results in significant dimension reduction. The minimal value of $d$ required for $\bm{\mathsf{\Gamma}}$ to satisfy (21) depends on the type of embedding $\bm{\mathsf{\Gamma}}$, the rank $k$ of $\bm{\mathsf{A}}$, and the distortion parameter $\epsilon$. As a classical example, the Johnson-Lindenstrauss lemma jlind84 establishes that $d\geq 8\log k/\varepsilon^{2}$ is sufficient for a random matrix $\bm{\mathsf{\Gamma}}$ with standard Gaussian entries to embed $k$ points in m so that (21) is satisfied.
More generally, if $\mathcal{S}$ is any probability distribution over linear maps $\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$, then $\bm{\mathsf{\Gamma}}\sim\mathcal{S}$ is a randomized embedding if (21) holds over $\mathcal{S}$ for all $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ with at least constant probability. If a randomized embedding can be constructed to map an unknown subspace into a lower-dimensional space (i.e. without knowing anything about the geometry of the input space except its dimension), then it is said to be an oblivious subspace embedding sarlos2006improved. It is often the case in practice that randomized DRMs are assumed to be randomized embeddings, but we will see in Sect. 3.1 that satisfying (21) is sufficient but not necessary to ensure the smaller-dimensional space is a good proxy for the input space.
In Sect. 3.1, we review randomized sketching, a procedure that produces random linear combinations of input matrix coordinates, which we say is “coordinate-mixing.” We then focus on dimension reduction by randomized sampling of matrix coordinates in Sect. 3.2. We end this section with an important application of randomized dimension reduction that will be relevant for our discussion of tensor decompositions: solving overdetermined least squares problems. For detailed treatments of randomized dimension reduction, the reader is referred to Camano2025_OSI; halko2011; martinsson2019randomized; martinsson2020; woodruff2015.
@@ 3.1 Randomized Matrix Sketching
The application of a randomized dimension reduction map $\bm{\mathsf{\Gamma}}$ to $\bm{\mathsf{A}}$ is known as randomized sketching. We reserve the term randomized sketching for random linear maps that are coordinate-mixing, vs. coordinate-sampling. The matrix product $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ is said to be a row sketch of $\bm{\mathsf{A}}$, and a column sketch $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is defined analogously.
Ideally, the random matrix used for sketching should roughly preserve the geometry of the row or column space of $\bm{\mathsf{A}}$. Randomized dimension reduction maps $\bm{\mathsf{\Gamma}}$ that are frequently utilized for sketching include:
  • Gaussian matrices ((halko2011, Section 4.1), (martinsson2020, Section 8.3), (woodruff2015, Theorem 2.3)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ with entries
    $\bm{\mathsf{\Gamma}}(i,j)\sim\mathcal{N}(0,1/d),$
    i.i.d. Gaussian entries drawn from a normal distribution with mean 0 and standard deviation $1/d$.
@@
    $\bm{\mathsf{\Gamma}}=\sqrt{\frac{m}{d}}\bm{\mathsf{\Pi}}_{m}\bm{\mathsf{F}}\bm{\mathsf{\Phi}}\bm{\mathsf{\Pi}}_{m\rightarrow d}.$
    Here, $\bm{\mathsf{\Pi}}_{m\rightarrow d}\in\mathbb{F}^{\hskip 0.56905ptm\times d}$ is a uniformly random selection of $d$ out of $m$ rows, whereas $\bm{\mathsf{\Pi}}_{m}\in\mathbb{F}^{\hskip 0.56905ptn\times n}$ is a random permutation of the $m$ rows. The matrix $\bm{\mathsf{F}}\in\mathbb{F}^{\hskip 0.56905ptm\times m}$ is a discrete Fourier transform, and $\bm{\mathsf{\Phi}}:=\textup{diag}(\phi_{1},\ldots,\phi_{n})$ has i.i.d. Rademacher random variable entries, i.e. when $\mathbb{F}=$, entries of $\pm 1$ chosen independently with equal probability.
  • Sparse sign matrices ((martinsson2020, Section 9.2), nelson2013, (woodruff2015, Theorem 2.5)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
    $\bm{\mathsf{\Gamma}}=\sqrt{\frac{m}{\zeta}}\begin{bmatrix}{\bm{\mathsf{s}}_{1},\ldots,\bm{\mathsf{s}}_{m}}\end{bmatrix}$
    for sparsity parameter $2\leq\zeta\leq d$, where each column $\bm{\mathsf{s}}_{j}\in\mathbb{F}^{\hskip 0.56905ptd}$ for $j=1,\ldots,m$ is filled with $\zeta$ independent Rademacher random variables at uniformly random coordinates. In other words, each column is an independent random vector having $\zeta$ nonzero entries of $\pm 1$ (when $\mathbb{F}=$) in uniformly random positions.
  • CountSketch matrices (Charikar2004; meng2013low; nelson2013): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as a sparse sign matrix with sparsity parameter $\zeta=1$. We caution that CountSketch matrices on their own can perform poorly in practice among randomized DRMs (see the required sketching dimension $d$ in Table 1).
@@
  • SparseCol matrices (Camano2025_OSI): For consistency with the previous definitions of row sketching matrices as subspace embeddings, we give the transpose of the SparseCol matrix in (Camano2025_OSI, Definition 4.1). Namely, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ is defined by its $d$ rows of sparse random vectors
@@
    where $\zeta$ is the sparsity parameter, $\rho_{1},\ldots\rho_{\zeta}$ are i.i.d. Rademacher random variables, the indices $s_{1},\ldots,s_{\zeta}$ are sampled uniformly from $[m]$ without replacement, and $\bm{\mathsf{e}}_{s_{i}}$ is the $s_{i}$-th standard basis vector. In other words, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ is the transpose of the SparseCol matrix $\bm{\mathsf{\Gamma}}^{*}\in\mathbb{F}^{m\times d}$ given by
    $\bm{\mathsf{\Gamma}}^{*}:=\frac{1}{\sqrt{k}}\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\ldots&\bm{\mathsf{\omega}}_{d}\end{bmatrix}$, (23)
    where $\bm{\mathsf{\omega}}_{j}\sim\bm{\mathsf{\omega}}$ are i.i.d.

### v2

@@ 2.2 The Singular Value Decomposition
where $r=\min(m,n)$, $\bm{\mathsf{U}}$ and $\bm{\mathsf{V}}$ are orthonormal, and $\mathbf{\Sigma}$ is diagonal. The columns $\{\bm{\mathsf{u}}_{i}\}_{i=1}^{r}$ and $\{\bm{\mathsf{v}}_{i}\}_{i=1}^{r}$ of $\bm{\mathsf{U}}$ and $\bm{\mathsf{V}}$, respectively, are called the left and right singular vectors of $\bm{\mathsf{A}}$. The diagonal elements $\{\sigma_{i}\}_{i=1}^{r}$ of $\mathbf{\Sigma}$ are the singular values of $\bm{\mathsf{A}}$, ordered so that $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$. The rank of $\bm{\mathsf{A}}$ is the number of nonzero singular values.
To obtain a rank-$k$ approximation of $\bm{\mathsf{A}}$, we can truncate its SVD after the first $k$ terms, defining $\bm{\mathsf{A}}_{k}=\sum_{i=1}^{k}\sigma_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{v}}_{i}^{*}$. By the Eckart-Young theorem eckart1936, $\bm{\mathsf{A}}_{k}$ is the best possible rank-$k$ approximation of $\bm{\mathsf{A}}$, with approximation error given by the singular values:
@@
We adopt this notation when referring to the computation of the SVD, as well as $\texttt{svd}(\bm{\mathsf{A}},k)$ to represent the function outputting the truncation $\bm{\mathsf{A}}_{k}$.
@@ 2.3 The QR Decomposition
Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits a QR decomposition $\bm{\mathsf{A}}=\bm{\mathsf{Q}}\bm{\mathsf{R}}$ where $\bm{\mathsf{Q}}$ is an orthonormal matrix of size $m\times r$, and where $\bm{\mathsf{R}}$ is an $r\times n$ upper triangular matrix, where $r=\min(m,n)$. In what follows, we are particularly interested in column pivoted QR factorizations, which take the form
$\underset{m\times n}{\hskip 5.69054pt\bm{\mathsf{A}}\hskip 5.69054pt}\underset{n\times n}{\bm{\mathsf{P}}}=\underset{m\times r}{\bm{\mathsf{Q}}}\underset{r\times n}{\hskip 5.69054pt\bm{\mathsf{R}}\hskip 5.69054pt},$ (3)
for some permutation matrix $\bm{\mathsf{P}}$. The point here is that it is often possible to pick a permutation matrix $\bm{\mathsf{P}}$ in such a way that the truncation of (3) to the dominant terms forms a good low rank approximation to $\bm{\mathsf{A}}$. Common algorithms such as Golub-Businger with standard column pivoting perform well in practice, and more advanced pivoting techniques come with provable guarantees gu1996, cf. Sect 2.5. These methods are greedy, and can be halted after a fixed number $k$ steps. We use the notation
$\begin{bmatrix}{\bm{\mathsf{Q}},\bm{\mathsf{R}},\bm{\mathsf{P}}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \begin{bmatrix}{\bm{\mathsf{Q}}_{k},\bm{\mathsf{R}}_{k},\bm{\mathsf{P}}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}},k),$ (4)
to refer to functions for computing either a full or a partial QR factorization. It is sometimes convenient to store the permutation using an index vector $I$ such that $\bm{\mathsf{P}}=\bm{\mathsf{I}}(\colon,I)$, where $\bm{\mathsf{I}}$ is the identity matrix. If we need only the factor $\bm{\mathsf{Q}}$, whose columns form an orthonormal basis for the range of $\bm{\mathsf{A}}$, we use the notation
@@
2.4 The LU Decomposition
Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits an LU decomposition given by
$\underset{m\times m}{\bm{\mathsf{P}}}\underset{m\times n}{\hskip 5.69054pt\bm{\mathsf{A}}\hskip 5.69054pt}=\underset{m\times r}{\bm{\mathsf{L}}}\underset{r\times n}{\hskip 5.69054pt\bm{\mathsf{U}}\hskip 5.69054pt},$ (6)
where $r=\min(m,n)$, $\bm{\mathsf{P}}$ is a permutation matrix, $\bm{\mathsf{L}}$ is lower triangular and $\bm{\mathsf{R}}$ is upper triangular. We can halt the factorization after computation of the first $k$ terms to obtain a “partial” LU factorization $\bm{\mathsf{A}}\approx\bm{\mathsf{L}}_{k}\bm{\mathsf{U}}_{k}$.
$\begin{bmatrix}{\bm{\mathsf{L}},\bm{\mathsf{U}},P}\end{bmatrix}=\texttt{lu}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \begin{bmatrix}{\bm{\mathsf{L}}_{k},\bm{\mathsf{U}}_{k},I}\end{bmatrix}=\texttt{lu}(\bm{\mathsf{A}},k),$ (7)
denote functions that yield either the full or partial LU factorization, respectively, where $\bm{\mathsf{P}}$ is stored as a permutation vector $P$ and $I=P(1:k)$.
@@ 2.5 The Interpolative and CUR Decompositions
where $\bm{\mathsf{R}}=\bm{\mathsf{A}}(I,:)$ for $k$ linearly independent rows indexed by $I$; $\bm{\mathsf{C}}=\bm{\mathsf{A}}(:,J)$ for $k$ linearly independent columns indexed by $J$; and $\bm{\mathsf{U}}=\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$, though in practice, $\bm{\mathsf{U}}$ is often approximated by $\bm{\mathsf{A}}(I,J)^{\dagger}$. The columns of $\bm{\mathsf{C}}$ or rows of $\bm{\mathsf{R}}$ are known as skeletons.
@@
where the coefficient $\sqrt{k+1}$ is optimal and cannot be improved; here, $\bm{\mathsf{A}}_{k}$ is the optimal rank-$k$ approximation given by the truncated SVD. In practice, greedy pivoting algorithms, based on QR, LU, or the SVD, are frequently used to choose skeletons for ID and CUR decompositions. We discuss several in the following sections.
ID/CUR from QR with Column Pivoting (CPQR)
Assuming $m\geq n$, let $\bm{\mathsf{A}}^{(t)}$ represent the resulting matrix after $t$ steps of pivoting and updating for $t=0,\ldots,n-2$ (with $\bm{\mathsf{A}}^{(0)}=\bm{\mathsf{A}}$). At the $(t+1)$-th step, CPQR chooses the column pivot $j_{t+1}$ as the column of maximal Euclidean norm in the active submatrix,
@@
In practice, CPQR chooses pivots in an order that reveals the rank of $\bm{\mathsf{A}}$, so that the chosen skeletons form a good basis for its row or column space. While there are adversarial cases for which CPQR can fail to be rank-revealing (e.g., Kahan-type matrices kahan1966), these are rarely encountered; cf. trefethen1990. There are variations of CPQR with better theoretical guarantees, such as rank-revealing QR Chan1987; Ilse1994 or strong rank-revealing QR GuEis1996, but they are more costly to implement.
ID/CUR from LU with Partial Pivoting (LUPP)
For an ID or CUR decomposition using LUPP, row pivots are selected from the active submatrix of $\bm{\mathsf{A}}$, which is updated after each pivot selection via Schur complements (e.g., (golub2013, Algorithm 3.2.1)). Let $\bm{\mathsf{A}}^{(t)}$ be the resulting matrix after $t$ steps of pivoting and updating for $t=0,1,\ldots,n-2$. At the ($t+1$)-st step of LUPP, the largest-magnitude column entry of the active submatrix is selected as the next pivot entry:
@@
where $\bm{\mathsf{A}}^{(t)}(t+1,t+1)$ has been updated with $\bm{\mathsf{A}}^{(t)}(i_{t+1},t+1)$ following row permutation. If expressed as the output of the function $\begin{bmatrix}\bm{\mathsf{L}},\bm{\mathsf{U}},P\end{bmatrix}=\texttt{lu}(\bm{\mathsf{A}})$, the row skeletons at the $(t+1)$-th step are given by $I=P(1:t+1)$.
This algorithm leads to an exponential upper bound on the entries of $\bm{\mathsf{U}}$, which is tight for certain matrices (e.g., Kahan-type matrices kahan1966). Thus, LUPP can fail to be rank-revealing, but again this is rare in practice trefethen1990. There are more sophisticated pivoting strategies to lend rank-revealing properties to the LU factorization Chan84; Pan2000, but at significantly greater computational cost. We will see that randomization can equip LUPP with rank-revealing properties comparable to CPQR even in adversarial cases.
ID/CUR from a Single Truncated SVD
@@
By (16) and the upper triangularity of $\widehat{\bm{\mathsf{V}}}^{*}$,
@@
Adding up over all previous steps, we have
@@
Letting $i=k$ in (20), we achieve the desired bound in (13) in $O(mnk)$ operations.
@@ 3 Randomized Dimension Reduction: Sketching and Sampling
Randomized DRMs are often associated with randomized embeddings, linear maps drawn from random distributions that satisfy the following condition with high probability. Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be a matrix of rank $k\leq\min(m,n)$ and $\varepsilon\in(0,1)$ a distortion parameter. A linear map $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an embedding of $\bm{\mathsf{A}}$ with distortion $\varepsilon$ if
@@
Usually, $k\leq d\ll m$. The minimal value of $d$ required for $\bm{\mathsf{\Gamma}}$ to satisfy (21) depends on the type of embedding $\bm{\mathsf{\Gamma}}$, the rank $k$ of $\bm{\mathsf{A}}$, and the distortion parameter $\epsilon$. As a classical example, the Johnson-Lindenstrauss lemma jlind84 establishes that $d\geq 8\log k/\varepsilon^{2}$ is sufficient for $\bm{\mathsf{\Gamma}}$ with standard Gaussian entries to embed $k$ points in m to satisfy (21).
More generally, if $\mathcal{S}$ is any probability distribution over linear maps $\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$, then $\bm{\mathsf{\Gamma}}\sim\mathcal{S}$ is a randomized embedding if (21) holds over $\mathcal{S}$ for all $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ with at least constant probability. If a randomized embedding can be constructed to map an unknown subspace into a lower-dimensional space (i.e. without knowing anything about the input space except its dimension), then it is said to be an oblivious subspace embedding (OSE) sarlos2006improved. Many times in practice, randomized DRMs are treated as embeddings, but we will see in Sect. 3.1 that satisfying (21) is sufficient but not necessary for the smaller-dimensional space to be a good proxy for the input.
In Sect. 3.1, we review randomized sketching, a procedure that produces random linear combinations of input matrix coordinates, which we say is “coordinate-mixing.” We then focus on randomized sampling of matrix coordinates in Sect. 3.2. We end with an important application of randomized dimension reduction that will be relevant for our discussion of tensor decompositions: solving overdetermined least squares problems. For detailed treatments, the reader is referred to Camano2025_OSI; halko2011; martinsson2019randomized; martinsson2020; woodruff2015.
@@ 3.1 Randomized Matrix Sketching
The application of a randomized DRM $\bm{\mathsf{\Gamma}}$ to $\bm{\mathsf{A}}$ is known as randomized sketching. We refer to $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ as a row sketch of $\bm{\mathsf{A}}$, with a column sketch $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ defined analogously. We generally reserve the term “sketching” for random linear maps that are coordinate-mixing, as opposed to methods that simply sub-sample columns.
Ideally, the random matrix used for sketching should roughly preserve the geometry of the row or column space of $\bm{\mathsf{A}}$. Commonly used randomized DRMs $\bm{\mathsf{\Gamma}}$ include:
  • Gaussian matrices ((halko2011, Section 4.1), (martinsson2020, Section 8.3), (woodruff2015, Theorem 2.3)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ with entries $\bm{\mathsf{\Gamma}}(i,j)\sim\mathcal{N}(0,1/d).$
@@
    $\bm{\mathsf{\Gamma}}=\sqrt{m/d}\ \bm{\mathsf{\Pi}}_{m}\bm{\mathsf{F}}\bm{\mathsf{\Phi}}\bm{\mathsf{\Pi}}_{m\rightarrow d}.$
    Here, $\bm{\mathsf{\Pi}}_{m\rightarrow d}\in\mathbb{F}^{\hskip 0.56905ptm\times d}$ is a uniformly random selection of $d$ out of $m$ rows, whereas $\bm{\mathsf{\Pi}}_{m}\in\mathbb{F}^{\hskip 0.56905ptn\times n}$ is a random permutation of the $m$ rows. $\bm{\mathsf{F}}\in\mathbb{F}^{\hskip 0.56905ptm\times m}$ is a discrete Fourier transform, and $\bm{\mathsf{\Phi}}:=\textup{diag}(\phi_{1},\ldots,\phi_{n})$ has i.i.d. Rademacher entries.
  • Sparse sign matrices ((martinsson2020, Section 9.2), nelson2013, woodruff2015): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
    $\bm{\mathsf{\Gamma}}=\sqrt{m/\zeta}\begin{bmatrix}{\bm{\mathsf{s}}_{1},\ldots,\bm{\mathsf{s}}_{m}}\end{bmatrix}$
    for sparsity parameter $2\leq\zeta\leq d$, where each column $\bm{\mathsf{s}}_{j}\in\mathbb{F}^{\hskip 0.56905ptd}$, $j=1,\ldots,m$ contains $\zeta$ i.i.d. Rademacher random variables at uniformly random coordinates.
  • CountSketch matrices (Charikar2004; meng2013low; nelson2013): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as a sparse sign matrix with sparsity $\zeta=1$. We advise against using CountSketch matrices on their own.
@@
  • SparseCol matrices (Camano2025_OSI): For consistency with the previous definitions of row sketching matrices as subspace embeddings, we give the transpose of the SparseCol matrix in (Camano2025_OSI, Definition 4.1). Namely, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ is defined by its $d$ rows given by
@@
    where $\zeta$ is the sparsity parameter, $\rho_{1},\ldots\rho_{\zeta}$ are i.i.d. Rademacher random variables, the indices $s_{1},\ldots,s_{\zeta}$ are sampled uniformly from $[m]$ without replacement, and $\bm{\mathsf{e}}_{s_{i}}$ is the $s_{i}$-th standard basis vector. In other words, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ is the transpose of the SparseCol matrix $\bm{\mathsf{\Gamma}}^{*}\in\mathbb{F}^{m\times d}$ given by $\bm{\mathsf{\Gamma}}^{*}:=\frac{1}{\sqrt{k}}\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\ldots&\bm{\mathsf{\omega}}_{d}\end{bmatrix}$, $\bm{\mathsf{\omega}}_{j}\sim\bm{\mathsf{\omega}}$ i.i.d.